OpenStudy (anonymous):

An automobile accelerates from rest at 2.5 m/s 2 for 29 s. The speed is then held constant for 28 s, after which there is an acceleration of −3.8 m/s 2 until the automobile stops. What total distance was traveled? Answer in units of km

OpenStudy (anonymous):

hero can you help him

OpenStudy (anonymous):

agent 47?

OpenStudy (agent47):

V_i=0 a=2.5(m/s^2) t=29s V_f - ? this is the first part of the solution.

OpenStudy (anonymous):

okay

OpenStudy (agent47):

V_f = V_i+a*t=0+2.5*29=72.5(m/s)

OpenStudy (anonymous):

\[x=x_o+v_ot+\frac{1}{2}at^2\]

OpenStudy (agent47):

Final velocity is 72.5(m/s) <- this is before he decelerates. @CliffSedge that formula wouldn't really work here.

OpenStudy (agent47):

So before he decelerates, the distance he has traveled is: well actually @CliffSedge it would, lol sorry :)

OpenStudy (anonymous):

lol

OpenStudy (agent47):

d = V_i * t + (1/2)a*t^2 = 0 + (1/2)*2.5*29^2 = 1051.25m so before he starts decelerating he has traveled 1051.25 meters

OpenStudy (anonymous):

@Agent47 , yes it does. i just used it.

OpenStudy (agent47):

Now once he starts decelerating: V_i = 72.5 (m/s) <- he was travelling with this velocity before he started slowing down V_f = 0 (m/s) <- the car comes to a stop t = 28s a = -3.8(m/s) Again, use the same formula...

OpenStudy (anonymous):

It's just for the first part, though.

OpenStudy (agent47):

d = V_i * t + (1/2)a*t^2 = 72.5*28 + (1/2)*(-3.8)*29^2 = 432.1m

OpenStudy (agent47):

So in total, he has traveled: 1051.25+432.1 = 1483.35m now convert that..

OpenStudy (anonymous):

The second part uses it (or a reduced form of it because of constant velocity) and v=a*t. The third part uses that information and \[\Delta v^2=2a \Delta x\]

OpenStudy (agent47):

to do so, just divide it by 1000, there are 1000 meters in a kilometer, so the final answer is: 1483.35/1000= 1.48335km

OpenStudy (agent47):

or about 1.5km

OpenStudy (agent47):

woah waitttt

OpenStudy (agent47):

The speed is then held constant for 28 s... my second part of the soln is wrong.

OpenStudy (agent47):

d = V_i * t + (1/2)a*t^2 = 0 + (1/2)*2.5*29^2 = 1051.25m so before he starts decelerating he has traveled 1051.25 meters, and his speed is: V_f = V_i+a*t=0+2.5*29=72.5(m/s) Then he sticks with that speed for 28s, so he travels another: 28*72.5=28*72.5=2030m

OpenStudy (agent47):

And then it decelerates at -3.8m/s^2 until it comes to a stop, so, again: V_i=72.5 V_f=0 a=-3.8 t - ? (I want to know what time it took him to stop) a = (V_f-V_i)/t -3.8=(0-72.5)/t t=72.5/3.8=19.0789474s

OpenStudy (agent47):

and again this formula: d = V_i * t + (1/2)a*t^2 saves the day, so during the deceleration, he has traveled: d = V_i * t + (1/2)a*t^2 = 72.5*19.1+(1/2)*(-3.8)*(19.1)^2= 691.6m

OpenStudy (agent47):

Ok, now add them all up: 1051.25+2030+691.6 = 3772.85m, or 3.8km

OpenStudy (anonymous):

okay

OpenStudy (agent47):

So basically you split questions like this one into three parts: 1) accelerating 2) constant 3) decelerating

OpenStudy (anonymous):

can you help me with another problem